OpenStudy (anonymous):

quadratic equations : solve: y^2 = 10y ?

OpenStudy (anonymous):

\[y^2-10y=0\] \[y(y-10)=0\] is a start

OpenStudy (anonymous):

y^2 - 10y = 0 y (y-10) = 0 y = 0, 10